# Enhanced SPS Velocity-Adaptive Scheme: Access Fairness in 5G NR V2I Networks

### v1

@@ I Introduction
With the 3rd Generation Partnership Project (3GPP) Release 16, the initial Vehicle-to-Everything (V2X) standard grounded in the 5G New Radio (NR) was introduced as a supplement to Long-Term Evolution (LTE) V2X communication[1]. NR V2X supports two types of communication: mode-1 (centralized) and mode-2 (distributed)[2, 3]. Vehicles may operate within network coverage in mode-1, and resources are scheduled through the Base station(BS)[4, 5]. In contrast, mode-2 allows vehicles to autonomously allocate resources with the sensing-based semi-persistent scheduling (SPS) mechanism, significantly enhancing the flexibility of resource scheduling for vehicles[6]. Modern vehicles are often outfitted with cameras, LiDARs, and other sensors to sense the environment[7]. However, due to limited onboard computational resources, it is challenging for vehicles to process the large-scale data they produce[8]. To extract useful information, vehicles often rely on cloud computing by offloading complex computational tasks to a central cloud , which processes the data and returns the results[9]. However, traditional cloud computing is unsuitable for high-speed vehicular scenarios, as long-distance transmission introduces significant communication delays, which can be dangerous in mobile vehicular environments[10, 11]. To address this, Vehicle-to-Infrastructure (V2I) technology is applied in vehicular network scenarios to obtain real-time data that assists with driving[12]. In NR V2I mode-2, vehicles typically employ sensing-based SPS mechanisms for resource scheduling[13]. However, vehicles on different lanes often travel at varying speeds, resulting in different durations of time spent within an RSU’s coverage area[14]. Furthermore, transmission failures may occur, leading to an unequal amount of successfully transmitted data for vehicles of different speeds[15]. This difference, referred to as unfair access, often results in high-speed vehicles receiving less information than low-speed vehicles, which increases the likelihood of incorrect decisions and safety risks[16].
@@
The key contributions of this work are summarized as below [^1] Source code can be found at https://github.com/qiongwu86/Enhanced-SPS-Velocity-adaptive-Scheme-Access-Fairness-in-5G-NR-V2I-Networks :
@@
  • 3)
    We adopt the Non-dominated Sorting Genetic Algorithms-II (NSGA-II)[17] to solve the proposed multi-objective optimization problem. Experimental results validate the efficiency of the proposed method.
The structure of the rest of this work is outlined below: Section II provides an overview of related studies. Section III presents the system model. Section IV presents the fairness metric. Section V formulates the multi-objective optimization problem and describes its solution using the NSGA-II. Section VI details the experimental results. Section VII concludes the paper.
@@ II Related Work
In this section, we will review the related works relevant to this paper.
@@ II-A Sensing Based Semi-Persistent Scheduling
Some studies have been worked on the SPS within NR V2X Mode-2. In [18, 19], Daw et al. proposed a priority-based SPS scheme that categorizes emergency vehicles separately and introduced a complementary probabilistic collision mitigation mechanism to minimize the collision probability for high-priority vehicles in the network. In [20], Jeon et al. significantly reduced data packet conflicts in the C-V2X mode-4 by minimizing the uncertainty in resource selection for SPS. Specifically, they utilized a “lookahead” approach to eliminate collisions caused by unawareness of other users’ decisions.
@@ II-B Fairness of network
Several studies have proposed solutions to address the fair access of vehicle caused by varying vehicle speeds. In [21], Wan et al. proposed modifying the contention window in the IEEE 802.11p protocol to address vehicle access fairness, while jointly considering the network-wide age of information to minimize latency. In [22], Harigovindan et al. employed proportional fairness (PF) in multi-rate, multi-lane V2I networks, improving the overall data transfer efficiency. In [23], Wu et al. addressed vehicle access fairness in a platooning scenario by dynamically tuning the minimum contention window according to vehicle speeds under the IEEE 802.11p.
However, these studies have not considered the fairness issue under the NR V2I Mode-2 SPS. Due to the few research focusing on vehicle access fairness in 5G NR V2X Mode 2 SPS, this motivates us to undertake this work.
@@ III System Model
This section focuses on a detailed analysis of the system model.
@@ III-A Scene Model
As depicted in Fig.1, we consider a highway environment featuring $N$ lanes, with an RSU deployed along the roadside as an edge server. The lanes are divided into two directions, and vehicles arrive within the RSU’s communication range following a Poisson process. Once a vehicle enters the RSU’s communication range, it will undertake information transmission with the RSU and extract useful information. For the same direction, the speed of vehicles in the same lane is the same.
@@ III-B Sensing Based Semi-Persistent Scheduling
Each vehicle adopts 5G NR V2X mode 2 for data transmission and undergoes resource allocation through the SPS[24]. Specifically, as Fig.2, the channel is partitioned into subframes temporally, each being 1ms[25]. Within the frequency spectrum, the channel is decomposed into subchannels, each composed of multiple consecutive resource blocks (RBs). A subchannel can be divided into two parts: the side link control information (SCI)[26], occupying two RBs, and the transmission block (TB), occupying the remainder. One subchannel combined with one subframe is referred to as a physical resource block (PRB). Upon reserving a PRB, it generates a reselection counter (RC) randomly and decrements it by 1 after each transmission[27]. When $RC=0$, it will reselect resources with a probability of $1-P$; otherwise, it will continue to utilize the previous resources[28].
Vehicles autonomously select subchannels and employ the selected subchannels for transmitting TB. Simultaneously, according to the resource reservation interval (RRI)[29], they announce the subchannel chosen for transmitting the next TB to avoid other vehicles choosing the same subchannel. During the reselection process, vehicle initially identifies resources within the selection window,[30] where the selection window’s size $w$ is determined independently by the vehicle. Subsequently, the vehicle determines the resources that should be excluded, namely, according to the information provided by the perception window of size 1000ms[31]. The exclusion criteria are as follows:
  • 1)
    PRBs that have already been reserved by other vehicles will be excluded.
  • 2)
    PRBs with an average RSRP over time exceeding the threshold will be excluded.
After eliminating all the PRBs that do not satisfy the conditions, the vehicle randomly selects the remaining PRBs for transmission resources.
@@ IV-A Transmission rate
where, $R$ represents the coverage range of the BS, and ${v}_{i}$ denotes the speed of vehicle. Considering the limits of our proposed scheme, and based on Shannon’s theorem, ${C}_{i}$ can be expressed as:
@@
where ${P}_{o}^{i}$ represents the position of vehicle $i$, and ${P}_{o}^{B}$ represents the position of the BS. The ${P}_{o}^{i}$ can be described as:
@@
where ${\Lambda}_{0}$ represents the wave length, and $\cos\theta$ is the angle between the communication direction and vehicle’s direction of motion .
@@ IV-B Successful decoding probability
where ${{\delta}^{j}}_{COL}$ represents the data packet collisions’ probability between vehicle $i$ and vehicle $j$. Multiple vehicles can occupy the same Physical Resource Block (PRB). When several vehicles nearly simultaneously attempt to select resources, there is a possibility that they may choose the same PRB. Based on the model in [33], ${{\delta}^{j}}_{COL}$ can be expressed as:
@@
where ${P}_{O}$ represents the probability of overlap between the selection windows of vehicle $i$ and vehicle $j$, and ${P}_{SH|O}$ is the probability which vehicle $i$ and vehicle $j$ select resources from their shared selection window[34]. Assuming the overlap occurs, represents the average count of candidate PRBs shared between vehicle $i$ and vehicle $j$, and $N_{Ca}$ represents the average number of candidate PRBs[35]. The candidate PRBs are the PRBs that are available for transmission after the re-selection steps (2)/(3). The sensing window size is 1000ms[w219]. ${P}_{O}$ can be mathematically expressed as:
@@
where ${N}_{Sc}$ represents the number of subchannels, and ${N}_{Sh}$ is the shared resources’ number within the overlapping selection window. ${N}_{Sh}$ can be expressed as:
@@
${{\delta}^{j}}_{HD}$ represents the probability that vehicles use the same time slot for transmission. Due to the half-duplex nature of vehicle communication[36], the receiver cannot decode the data packet from the transmitter if both vehicles transmit simultaneously, leading to a transmission failure. Based on the model in [33], this can be expressed as:
@@ IV-C Fairness index
For brevity, we only consider the fairness index at a certain time. Based on the above analysis, Eq.(3) can be further expressed as:
@@
where,
${C}^{\prime}=B\cdot R\cdot\prod\limits_{j\neq\text{i}}{(1-{{\delta}^{j}}_{HD})}.$ (18)
Thus, we have derived the fairness index for vehicle $i$. Furthermore, since ${d}_{i}$ is a function of $v_{i}$ and ${{P}_{PRR}}^{{}_{i}}$ is a function of $w$, the fairness index $K_{index}^{i}$ is a function of both $v$ and $w$. Therefore, by knowing the vehicle’s speed, we can adaptively adjust the vehicle’s selection window based on speed to achieve overall network fairness.
@@
$K_{index}={{\log}_{2}}(1+\frac{{{p}_{i}}\cdot{{h}_{i}}\cdot{{({{d}_{i}}(\bar{v}))}^{-\partial}}}{{{\sigma}^{2}}})\cdot\frac{\prod\limits_{j\neq\text{i}}{(1-{{\delta}^{j}}_{COL}(\bar{w}))}}{{{{{v}}}_{i}}}$ (19)
where $\bar{v}$ represents the network’s average speed, and $\bar{w}$ represents the network’s average window size. The fairness index $K_{index}$ can measure the overall network’s fairness. The vehicle is achieving fair access when $K_{index}^{i}$ approaches $K_{index}$.
@@ V Optimization Problem and Solution
In this section, we formulate a multi-objective optimization problem and employ the NSGA-II algorithm [17]. As a result, we can derive the optimal selection window size that achieves network fairness.
@@ V-A Optimization Objective
The optimization goal is to adjust each vehicle’s selection window sizes to ensure that the data transmitted between each vehicle and the RSU is similar. This implies that the fairness index $K_{index}^{i}$ approaches the network’s average fairness index $K_{index}$. Accordingly, the optimization objective functions can be formulated as:
Objective 1 to $\bm{N}$: Minimize the variation between the fairness index on different lanes and network’s fairness index.
$\min_{\bm{w}}$ $\bm{F}(\bm{w})=[F_{K_{1}}(\bm{w}),F_{K_{2}}(\bm{w}),\dots,F_{K_{N}}(\bm{w})]^{T}$ (20)
@@
$F_{K_{i}}(\bm{w})=\left|K_{index}(\bm{w})-K_{index}^{i}(\bm{w})\right|,i\in[1,\ldots,N],$ (21)
$\bm{w}=\{w^{1},w^{2},...,w^{N}\}$. $w^{LB}$ and $w^{UB}$ indicate the minimum and maximum bounds of the selection window sizes, as defined in the NR V2I SPS based on [6].
Thus, we have defined a multi-objective optimization problem, which enables achieving overall network fairness by adjusting the selection window sizes of individual vehicles.
@@ V-B Optimization Solution
Input: $\bm{v}=\{v_{1},v_{2},\dots,v_{N}\}$, $N_{max}$: Maximum number of generations, $M$: population size, Threshold
Output: optimal solution $\bm{w}^{*}$
1 Initialize the population $P_{0}=\{\bm{w_{1}},\bm{w_{2}},\dots,\bm{w_{M}}\}$ within $[w^{LB},w^{UB}]$ randomly;
@@
3       Crossover: Generate offspring $P_{n}^{C}$ from $P_{n}$ using crossover;
@@ 5       Population Merge: Form $Q_{n}=P_{n}\cup P_{n}^{M}$;
6       Objective function value: Calculate the value of objective function $F_{K}(\bm{w_{i}})=\{f_{1}(\bm{w_{i}}),f_{2}(\bm{w_{i}}),\dots,f_{M}(\bm{w_{i}})\}$ based on Eq.(21), for all $\bm{w_{i}}\in Q_{n}$;
7       Non-dominated Sorting: Compute $\bm{F}=\{F_{1},F_{2},\dots\}$ for $Q_{n}$ with $F_{K}(\bm{w_{i}})$;
8       Crowding Distance: Compute crowding distance $d(\bm{w_{i}})$ for all $\bm{w_{i}}\in Q_{t}$ with $F_{K}(\bm{w_{i}})$;
9       Selection: Select the top $M$ individuals in $Q_{n}$ as $P_{n+1}$ using $\bm{F}$ and $d(\bm{w_{i}})$;

### v3

@@ I Introduction
With the 3rd Generation Partnership Project (3GPP) Release 16, the initial Vehicle-to-Everything (V2X) standard grounded in the 5G New Radio (NR) was introduced as a supplement to Long-Term Evolution (LTE) V2X communication. NR V2X supports two types of communication: mode-1 (centralized) and mode-2 (distributed) [1, 2]. Vehicles may operate within network coverage in mode-1, and resources are scheduled through the Base station(BS). In contrast, mode-2 allows vehicles to autonomously allocate resources with the sensing-based semi-persistent scheduling (SPS) mechanism, significantly enhancing the flexibility of resource scheduling for vehicles. Modern vehicles are often outfitted with cameras, LiDARs, and other sensors to sense the environment[3, 4]. However, due to limited onboard computational resources, it is challenging for vehicles to process the large-scale data they produce [5, 6]. To address this, Vehicle-to-Infrastructure (V2I) technology is applied in vehicular network scenarios to obtain real-time data that assists with driving[7, 8]. In NR V2I mode-2, vehicles typically employ sensing-based SPS mechanisms for resource scheduling[9]. However, vehicles on different lanes often travel at varying speeds, resulting in different durations of time spent within an RSU’s coverage area. Furthermore, transmission failures may occur, leading to an unequal amount of successfully transmitted data for vehicles of different speeds[10, 11]. This difference, referred to as unfair access, often results in high-speed vehicles receiving less information than low-speed vehicles, which increases the likelihood of incorrect decisions and safety risks.
@@
The contributions of this work are summarized :
@@
The structure of the rest of this work is outlined below: Section II provides an overview of related works. Section III presents the system model. Section IV presents the fairness index. Section V formulates the optimization problem and describes solution using NSGA-II. Section VI details the experimental results. Section VII concludes the paper.
@@ II Related Work
In this section, we first introduce the relevant improvements related to SPS, followed by a discussion on the enhancements made to network fairness.
@@ II-A Sensing Based Semi-Persistent Scheduling
Some studies have been worked on the SPS within NR V2X Mode-2. In [12], Daw et al. proposed a priority-based SPS scheme that categorizes emergency vehicles separately and introduced a complementary probabilistic collision mitigation mechanism to minimize the collision probability for high-priority vehicles in the network. In [13], Jeon et al. significantly reduced data packet conflicts in the C-V2X mode-4 by minimizing the uncertainty in resource selection. Specifically, they utilized a “lookahead” approach to eliminate collisions caused by unawareness of other users’ decisions. In [14], Dayal et al. extended SPS to accommodate adaptive RRI, referred to as SPS++, to address the severe underutilization and overutilization of radio resources in various vehicular traffic scenarios. This issue significantly impairs the timely dissemination of BSMs, thereby increasing the risk of collisions. In [15], Gu et al. proposed an SPS analytical model that quantifies the impact of beacon rate, range settings, and system configuration on the probability of access collisions and delay outage. The analytical model provides critical insights and guidance for adapting and optimizing protocol parameters, including sensing range, transmission power, and resource reservation.
The studies above have improved the SPS mechanism from various perspectives. However, none of these works have considered leveraging the sensing window size to address the issue of access fairness in vehicular networks.
@@ II-B Fairness of network
Several studies have proposed solutions to address the fair access of vehicle caused by varying vehicle speeds. In [16], Wan et al. proposed modifying the contention window in the IEEE 802.11p protocol to address vehicle access fairness. In [17], Wu et al. addressed vehicle access fairness in a platooning scenario by dynamically tuning the minimum contention window according to vehicle speeds under the IEEE 802.11p. In [18], Praghash et al. proposed a metric to achieve a certain level of fairness among network users and employed a reinforcement learning (RL) algorithm to mitigate conflicts between clients. In [19], Song et al. proposed a novel two-phase scheme, termed Energy-aware UAV Relay Transmission (EURT), where the users’ limited residual energy (RE) leads to unfair transmission delays within the network. The scheme aims to balance users’ transmission delays and network fairness.
However, the studies above have failed to simultaneously consider the issue of fair access in vehicular networks, particularly under the 5G NR V2X protocol. Only a few works have addressed fair access under the IEEE 802.11p protocol. Moreover, since the current SPS mechanism cannot flexibly adjust the selection window size based on vehicle speed, it fails to dynamically maintain V2I access fairness according to vehicle speed. Consequently, high-speed vehicles encounter difficulties in exchanging sufficient information with the RSU. Therefore, we are motivated to conduct this study.
@@ III-A Scene Model
As depicted in Fig.1, we consider a highway environment featuring $N$ lanes, with an RSU deployed along the roadside as an edge server. The lanes are divided into two directions. vehicles arrive within the RSU following a Poisson process. Once a vehicle enters the RSU’s communication range, it will undertake information transmission with the RSU and extract useful information.
@@ III-B Sensing Based Semi-Persistent Scheduling
Each vehicle adopts 5G NR V2X mode 2 for data transmission and undergoes resource allocation through the SPS.
Specifically, as Fig.2, the channel is partitioned into subframes, each being 1ms. Within the frequency spectrum, the channel is decomposed into subchannels, each composed of multiple consecutive resource blocks (RBs). A subchannel can be divided into two parts: the side link control information (SCI), occupying two RBs, and the transmission block (TB), occupying the remainder. One subchannel combined with one subframe is referred to as a physical resource block (PRB). Upon reserving a PRB, it generates a reselection counter (RC) randomly and decrements it by 1 after transmission. When $RC=0$, it will reselect resources with a probability of $1-P$, otherwise, it will continue to utilize the previous resources. During the reselection process, vehicle initially identifies resources within the selection window, where the selection window’s size $w$ is determined independently by the vehicle. Subsequently, the vehicle determines the resources that should be excluded, namely, according to the information provided by the perception window of size 1000ms.
@@ IV-A Transmission rate
where $R$ represents the coverage range of the RSU, and ${v}_{i}$ denotes the speed of vehicle. Based on Shannon’s theorem, ${C}_{i}$ can be expressed as:
@@
where ${P}_{o}^{i}$ represents the position of vehicle $i$, and ${P}_{o}^{B}$ is the position of the BS. The ${P}_{o}^{i}$ can be described as:
@@
where ${\Lambda}_{0}$ is the wave length, and $\cos\theta$ is the angle between the communication direction and the direction of motion.
@@ IV-B Successful decoding probability
where ${{\delta}^{j}}_{COL}$ represents the data packet collisions’ probability between vehicle $i$ and vehicle $j$. When several vehicles nearly simultaneously attempt to select resources, there is a possibility that they may choose the same PRB. Based on the model in [21], ${{\delta}^{j}}_{COL}$ can be expressed as:
@@
where ${P}_{O}$ is the probability of overlap between the selection windows of vehicle $i$ and vehicle $j$. ${P}_{SH|O}$ is the probability which vehicle $i$ and vehicle $j$ select resources from their shared selection window. $N_{Ca}$ is the average number of candidate PRBs. ${P}_{O}$ and ${P}_{SH|O}$ can be expressed as:
@@
where ${N}_{Sc}$ is the number of subchannels. ${N}_{Sh}$ is the shared resources number within the overlapped selection window.
${N}_{Sh}$ can be expressed as:
@@
${{\delta}^{j}}_{HD}$ represents the probability that vehicles use the same time slot for transmission. Due to the half-duplex nature of vehicle communication, the receiver cannot decode the data packet from the transmitter if both vehicles transmit simultaneously. Based on the model in [21], this can be expressed as:
@@ IV-C Fairness index
For brevity, we only consider the fairness index at a certain time. Eq.(3) can be further expressed as:
@@
Thus, we have derived the fairness index for vehicle $i$. Furthermore, since ${d}_{i}$ is a function of $v_{i}$ and ${{P}_{PRR}}^{{}_{i}}$ is a function of $w$, $K_{index}^{i}$ is a function of both $v$ and $w$.
Therefore, by knowing the vehicle’s speed, we can adaptively adjust the vehicle’s selection window based on speed to achieve overall network fairness.
@@
$K_{index}={{\log}_{2}}(1+\frac{{{p}_{i}}{{h}_{i}}{{{{d}_{i}}(\bar{v})}^{-\partial}}}{{{\sigma}^{2}}})\cdot\frac{\prod\limits_{j\neq\text{i}}{(1-{{\delta}^{j}}_{COL}(\bar{w}))}}{\bar{v}}$ (18)
where $\bar{v}$ represents the average speed, and $\bar{w}$ represents the network’s average window size. The fairness index $K_{index}$ can measure the overall network’s fairness. The vehicle is achieving fair access when $K_{index}^{i}$ approaches $K_{index}$.
@@ V Optimization Problem and Solution
In this section, we formulate a multi-objective optimization problem and employ the NSGA-II algorithm [22].
@@ V-A Optimization Objective
The optimization goal is to adjust vehicle’s selection window sizes to ensure that data transmitted between each vehicle and the RSU is similar. This implies that $K_{index}^{i}$ approaches the network’s fairness index $K_{index}$. Accordingly, the optimization objective functions can be formulated as:
Objective 1 to $\boldsymbol{N}$: Minimize the variation between the fairness index on different lanes and network’s fairness index.
$\min_{\boldsymbol{w}}$ $\boldsymbol{F}(\boldsymbol{w})=[F_{K_{1}}(\boldsymbol{w}),F_{K_{2}}(\boldsymbol{w}),\dots,F_{K_{N}}(\boldsymbol{w})]^{T}$ (19)
@@
$F_{K_{i}}(\boldsymbol{w})=\left|K_{index}(\boldsymbol{w})-K_{index}^{i}(\boldsymbol{w})\right|,$ (20)
$\boldsymbol{w}=\{w^{1},w^{2},...,w^{N}\}$. $w^{LB}$ and $w^{UB}$ indicate the minimum and maximum bounds of the selection window sizes. Our aim is to filter the resulting set of Pareto solutions to find the optimal solution.
@@ V-B Optimization Solution
We use the NSGA-II algorithm [22] to address the optimization problem. Each population consists of $N$ individuals, representing the vehicle selection windows. The population size is $M$. The algorithm is presented in Algorithm 1.
2) Iteration Phase: Firstly, crossover and mutation are performed. Crossover involves exchanging parts of two parent individuals’ genes to generate new offspring. Mutation randomly modifies an individual’s genes to prevent the algorithm from converging to a local optimum. At this point, the parent and offspring populations are merged into a new population $Q_{n}$ for the subsequent selection operation.
Next, we use $\boldsymbol{v}=\{v_{1},v_{2},\dots,v_{N}\}$ to calculate the objective function values $F_{K}(\boldsymbol{w})$ according to Eq.(20).
Then we perform non-dominated sorting. Non-dominated sorting divides the population into multiple levels according to each individual’s dominance rank. Thus we derive $\boldsymbol{F}=[F_{1},F_{2},\dots]$, $F_{1}$ means the individuals which rank first, and then, crowding distance is calculated. Parent individuals are selected based on non-dominated sorting and crowding distance. The selection prefers individuals that are ranked higher within the same rank. Individuals with larger crowding distances are preferred.
Thus, the merged population $Q_{n}$ undergoes non-dominated sorting and selection. Finally, $M$ individuals with higher domination rank and larger crowding distances are selected to form the new population $p_{n+1}$.
3) Optimization Phase: The optimal population is selected from the Pareto front. The optimal population must ensure that the difference between the fairness index of each vehicle is smaller than a threshold.
Meet the above conditions while also minimizing the total objective function values. Thus, the optimal selection window $\boldsymbol{w}^{*}$ is obtained.
Input: $\boldsymbol{v}=\{v_{1},v_{2},\dots,v_{N}\}$, $N_{max}$: Maximum number of generations
Output: optimal solution $\boldsymbol{w}^{*}$
1 Initialize the population $P_{0}=\{\boldsymbol{w_{1}},\boldsymbol{w_{2}},\dots,\boldsymbol{w_{M}}\}$ within $[w^{LB},w^{UB}]$ ;
@@
3       Crossover: Generate $P_{n}^{C}$ from $P_{n}$ with crossover;
@@ 5       Population Merge: Form $Q_{n}=P_{n}\cup P_{n}^{M}$;
6       Objective function value: Calculate $F_{K}(\boldsymbol{w_{i}})=\{f_{1}(\boldsymbol{w_{i}}),f_{2}(\boldsymbol{w_{i}}),\dots,f_{M}(\boldsymbol{w_{i}})\}$ based on Eq.(20), for all $\boldsymbol{w_{i}}\in Q_{n}$;
7       Non-dominated Sorting: Compute $\boldsymbol{F}=\{F_{1},F_{2},\dots\}$ for $Q_{n}$;
8       Crowding Distance: Compute crowding distance $d(\boldsymbol{w_{i}})$ for all $\boldsymbol{w_{i}}\in Q_{n}$ with $F_{K}(\boldsymbol{w_{i}})$;
9       Selection: Select the top $M$ individuals in $Q_{n}$ as $P_{n+1}$ using $\boldsymbol{F}$ and $d(\boldsymbol{w_{i}})$;